Calculate the pH of a carbonate buffer system that contains 0.0065 M HCO3- and 0.0084 M CO32-. ( Ka of HCO3- is 4.7x0-11 )

3Calculate the pH of a carbonate buffer system that contains 0.0065 M HCO


1
Expert's answer
2021-02-01T04:03:20-0500

PH = pKa - log(acid/salt)

= -log (4.7 × 10-11) - log (0.0065/0.0084)

= 10.33- (-0.11)

= 10.44